### v1

@@ 1 Introduction
Cannings models, first introduced in [cannings74, cannings1975], form a class of discrete-time population models in which each generation consists of a fixed number of individuals, and reproduction is governed by an exchangeable offspring distribution. This framework includes many classical models such as the Wright–Fisher model and (discrete-time) Moran model, and has played a central role in mathematical population genetics.
In this paper, we study the genealogical trees arising from Cannings models, referred to as Cannings trees. While the scaling limits of various genealogical trees—most notably Galton–Watson trees—have been extensively studied (e.g., [1993Aldous, marckert2003depth, le2006random]), the scaling behavior of Cannings trees remains largely unexplored. This motivates our present work.
To obtain a richer class of limiting results, we extend the classical Cannings framework to a class of inhomogeneous Cannings models, in which the population size may vary across generations. We assume that the rescaled generation-size profile converge to a continuous function $\ell$. Our main result (Theorem 1.1) shows that, under suitable moment assumptions, the rescaled height and contour functions of Cannings trees converge to a time change of Brownian motion conditioned to have local time profile $\ell$.
@@ 1.1 Setup and main results
Let $W$ be a reflected Brownian motion on $\mathbb{R}^{+}=[0,\infty)$, and denote by $L_{t}(x)$ its local time at level $x$ up to time $t$. Define the inverse local time at the origin by
$\tau_{u}:=\inf\{t\geq 0:L_{t}(0)>u\}$
A function $\ell:\mathbb{R}^{+}\to\mathbb{R}^{+}$ is called a continuous profile if it satisfies:
@@
This conditioned process was constructed explicitly by Warren–Yor [warren98] and characterized via continuum random trees by Aldous [1998BROWNIAN]. In Section 2.3, we recall these constructions and introduce a canonical version of $W^{\ell}$. Throughout the paper, we always work with that canonical version.
@@
      and
[…]
@@
      $\lim\limits_{s\rightarrow 0+}\dfrac{\ell(s)}{\sigma(s)^{2}}$ exists.
@@
    One of the following conditions holds:
    1. (i)
      For every $\delta>0$, there exists $\varepsilon>0$ such that uniformly for $\delta n\leq s\leq h(q_{n})-\delta n$,
      $\mathbb{E}\big[(\nu_{1}^{n,s})^{3}\big]=o\!\left(\frac{n}{(\log n)^{2+\varepsilon}}\right);$
    2. (ii)
      For every $\delta>0$, the family $\big\{\nu^{n,s}_{1}:n\geq 1,\ \delta n\leq s\leq h(q_{n})-\delta n\big\}$ is uniformly $L^{2}$-integrable, i.e.
      $\lim_{K\to\infty}\ \sup_{n\geq 1}\ \sup_{\delta n\leq s\leq h(q_{n})-\delta n}\mathbb{E}\big[(\nu_{1}^{n,s})^{2}\,\mathbf{1}\{\nu^{n,s}_{1}>K\}\big]=0.$ (1)
Remark. Condition H1(ii) in (H1) ensures that the contribution to the rescaled population size from the very early and very late generations—near the beginning and near extinction—is negligible. Combined with H1(i), this yields the convergence of the rescaled total population size:
$\frac{1}{n^{2}}\sum_{s\in\mathbb{N}_{+}}q_{n}(s)\;\longrightarrow\;\int_{0}^{\infty}\ell(s)\,\mathrm{d}s.$ (2)
@@
extends continuously to $s=0$, and hence defines a continuous profile. As will be seen in Theorem 1.1, the limiting object involves Brownian motion conditioned on the local time profile $\ell^{\sigma}$. Hence condition H1(iii) rules out the pathological case we would otherwise have to handle—conditioning on a profile whose value at the origin fails to converge.
@@
Assume (H1)–(H2). Define $\alpha(t):=\frac{1}{4}\int_{0}^{t}\sigma(W^{\ell^{\sigma}}_{s})^{2}\,\mathrm{d}s$. Let $\alpha^{-1}(t)$ be the inverse of $\alpha(t)$ [^4] We will verify in Lemma 2.16 that $\alpha$ is a.s. strictly increasing, so $\alpha^{-1}(t)$ is well-defined.. Then
$\Big(n^{-1}C_{\lfloor 2n^{2}t_{1}\rfloor}(\mathscr{T}^{n}),n^{-1}H_{\lfloor n^{2}t_{2}\rfloor}(\mathscr{T}^{n})\Big)_{0\leq t_{1},t_{2}<\infty}\overset{d}{\longrightarrow}\Big(W_{\alpha^{-1}(t_{1})}^{\ell^{\sigma}},W^{\ell^{\sigma}}_{\alpha^{-1}(t_{2})}\Big)_{0\leq t_{1},t_{2}<\infty}$ (3)
@@
It is a standard fact that convergence of contour functions implies convergence of the associated trees in the Gromov-Hausdorff topology (see [le2006random, Lemma 2.4]). Applying this, we obtain the following corollary.
@@
Assume (H1)–(H2). Let $\widetilde{\mathscr{T}}^{n}$ be the geometric tree obtained from $\mathscr{T}^{n}$ by assigning length $1/n$ to each edge. Let $\mathcal{T}^{\ell,\sigma}$ denote the continuum random tree encoded by the time-changed conditioned Brownian motion $W^{\ell^{\sigma}}_{\alpha^{-1}(t)}$. Then
@@
Condition (H2) is a sufficient but not optimal assumption for the convergence result (3). A sharper condition consists of (H2’) (stated in Section 3.3) together with
$\lim_{n\rightarrow\infty}\ \sup_{\delta n\leq s\leq h(q_{n})-\delta n}\frac{\mathbb{E}\big[(\nu_{1}^{n,s})^{3}\big]}{n}=0.$ (H2”)
The corresponding convergence result under (H2’) and (H2”) is given in Theorem 5.6. We show in Proposition 5.8 that the condition (H2) implies both (H2’) and (H2”).
By Aldous’ theorem [1993Aldous, Theorem 20], the convergence of the contour functions is equivalent to the combination of tightness and finite-dimensional convergence of the associated tree sequence. The condition (H2”) is in fact equivalent to the finite-dimensional convergence (see Remark 5.3); while condition (H2’) is closely related to the tightness. In Section B, we present a counterexample in which the offspring distribution slightly violates (H2’), leading to a failure of tightness. This indicates a form of sharpness of condition (H2’).
We also remark that the counterexample in Section B shows that the condition (H2”), although sufficient for finite-dimensional convergence, is not enough to guarantee tightness.
@@
The convergence results can also be extended to other population models, such as the continuous-time Moran model and the diploid Cannings model (as defined in [mohle2003coalescent]), using a very similar approach.
@@ 1.2 Related works
A major focus in the study of Cannings models lies in coalescent theory, which describes the genealogies of finite samples via limiting processes such as Kingman’s coalescent [kingman1982genealogy]. As explained in Remark 5.3, convergence to Kingman’s coalescent is in fact equivalent to the finite-dimensional convergence of the associated tree sequence.
In [mohle2000total, Theorem 4], Möhle gave the necessary and sufficient condition for convergence to Kingman’s coalescent in the classical Cannings model. Condition (H2”) serves as the analogue of this criterion in the inhomogeneous Cannings setting. Since it is equivalent to the finite-dimensional convergence, it constitutes a necessary condition of the convergence result (3), as discussed in Remark 1.3.
@@
It is classical that every continuous bridge function naturally encodes a continuum random tree (see Section 2.2.2 for details). For the conditioned Brownian motion $W^{\ell}$, Aldous [1998BROWNIAN] derived the finite-dimensional distributions of the corresponding encoding tree. Subsequently, Warren [warren1999result] provided two alternative, direct proofs of this result. This distribution plays a key role in identifying the limiting process in our setting.
@@
The conditioned Brownian motion $W^{\ell}$ is a self-interacting diffusion constructed by Warren and Yor [warren98], and independently by Aldous [1998BROWNIAN]. Detailed constructions will be presented in Section 2.3. Subsequently, Warren [warren05] further established a close connection between this conditioning problem and a stochastic flow of the Bass-Burdzy type.
More recently, Lupu, Sabot, and Tarrès [lupu2021inverting], as well as Aïdékon, Hu, and Shi [aidekon2023stochastic], have further developed this framework by analyzing perturbed reflecting Brownian motions and Brownian loop soups conditioned on local time profiles [^5] Specifically, Lupu, Sabot, and Tarrès [lupu2021inverting] studied the Gaussian free field conditioned on local time profile, which is equivalent—via Dynkin’s isomorphism—to a Brownian loop soup of intensity $1/2$, or equivalently, a perturbed reflecting Brownian motion with parameter $2$, conditioned on local time profile. Aïdékon, Hu, and Shi [aidekon2023stochastic] later generalized this result to arbitrary loop soup intensities (or equivalently, arbitrary parameters for perturbed reflecting Brownian motion).. A natural extension of our setting is to incorporate immigration into the inhomogeneous Cannings model. In that case, the rescaled contour and height processes will converge to the (time-changed) perturbed reflecting Brownian motion conditioned on its local time profile.
On the discrete side, the conditional laws of Markov chains and Markovian loop soups given their local time profiles have been studied in [2016Inverting, lupu2019inverting, li2024inverting].
@@ 1.3 Proof outline
Our approach to proving Theorem 1.1 builds on the framework developed by Aldous in [1993Aldous, Theorem 20], where he showed that for a sequence of random rooted ordered trees, convergence of the contour process is equivalent to the combination of finite-dimensional convergence and tightness of the trees (see Theorem 2.6 for a modified version of this result adapted to our setting). In our case, finite-dimensional convergence is obtained via its equivalence with convergence to Kingman’s coalescent (see Proposition 5.1). The main technical challenge lies in proving tightness. Once tightness is established, the convergence of the contour process follows, and the joint convergence with the height process is then deduced using a deterministic relation between the two processes from [marckert2003depth] (see Section 5.2).
@@
We highlight that the first fact relies on a sequential version of the coming-down-from-infinity property, established in Theorem 3.26. To this end, we introduce a coalescent process, a Markov process that tracks the number of ancestral lineages backward from height $(s+1)\lfloor\delta n\rfloor$. In particular, at height $s\lfloor\delta n\rfloor$ this count is exactly the number of $\delta$-coalescent vertices at the $s$-th height interval. Our assumption (H2’) is used in the proof to control the merge rates of these lineages. Heuristically, under (H2’) the process contracts at a uniform multiplicative rate on mesoscopic windows: on average, over each window of length $\frac{n}{(\log n)^{1+\varepsilon}},$ the lineage count shrinks by a factor $(\log n)^{-c}$. So over a time horizon $\delta n$, the cumulative shrinkage is
@@
This results in a rapid decrease in the number of lineages—from $\Theta(n)$ down to a constant—within $\delta n$ steps. By contrast, in the counterexample of Section B (where (H2’) fails), a contraction occurs only every $\frac{n}{(\log n)^{1-\varepsilon}}$ steps; over the same horizon $\delta n$ the cumulative shrinkage is
@@
which is insufficient to reduce the lineage count from order $\Theta(n)$ to $O(1)$. This leads to non-convergence of the rescaled contour functions in this example (see Proposition B.1).

### v2

@@
Abstract
We study the scaling limits of genealogical trees arising from Cannings models. Under suitable moment conditions, we show that the rescaled contour and height functions converge to a time change of Brownian motion conditioned on a given local time profile. This conditioned Brownian motion is a self-interacting diffusion constructed independently by Warren–Yor (1998) and Aldous (1998). A key ingredient in our proof is a sequential version of the coming-down-from-infinity property.
Abstract
[language=french] Nous étudions les limites d’échelle des arbres généalogiques issus de modèles de Cannings. Sous des conditions de moments appropriées, nous montrons que les fonctions de contour et de hauteur, après renormalisation, convergent vers un changement de temps d’un mouvement brownien conditionné par un profil de temps local donné. Ce mouvement brownien conditionné est une diffusion auto-interagissante, construite indépendamment par Warren–Yor (1998) et Aldous (1998). Un ingrédient clé de notre preuve est une version séquentielle de la propriété de descente depuis l’infini.
conditioned Brownian motion,
keywords:
[class=MSC]
keywords:
@@ 1 Introduction
Cannings models, first introduced in [5, 6], form a class of discrete-time population models in which each generation consists of a fixed number of individuals, and reproduction is governed by an exchangeable offspring distribution. This framework includes many classical models such as the Wright–Fisher model and (discrete-time) Moran model, and has played a central role in mathematical population genetics.
In this paper, we study the genealogical trees arising from Cannings models, referred to as Cannings trees. While the scaling limits of various genealogical trees—most notably Galton–Watson trees—have been extensively studied (e.g., [2, 13, 8]), the scaling behavior of Cannings trees remains largely unexplored. This motivates our present work.
To obtain a richer class of limiting results, we extend the classical Cannings framework to a class of inhomogeneous Cannings models, in which the population size may vary across generations. We assume that the rescaled generation-size profiles converge to a continuous function. Our main result (Theorem 1.1) shows that, under suitable moment assumptions, the rescaled height and contour functions of Cannings trees converge to a time change of Brownian motion conditioned on local time profile.
@@ 1.1 Setup and main results
Let $W$ be a reflected Brownian motion on $\mathbb{R}_{+}=[0,\infty)$, and denote by $L_{t}(x)$ its local time at level $x$ up to time $t$. Define the inverse local time at the origin by
$\tau_{u}:=\inf\{t\geq 0:L_{t}(0)>u\}.$
A function $\ell:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is called a continuous profile if it satisfies:
@@
This conditioned process was constructed explicitly by Warren–Yor [20] and characterized via continuum random trees by Aldous [3]. In Section 2.3, we recall these constructions and introduce a canonical version of $W^{\ell}$. Throughout the paper, we always work with that canonical version.
@@
      $\text{and}\quad\sigma_{n}(\lfloor n\,\cdot\rfloor)\to\sigma\ \text{uniformly on compact subsets of }(0,h(\sigma)).$
@@
      Both $\lim\limits_{s\rightarrow 0+}\dfrac{\ell(s)}{\sigma(s)^{2}}$ and $\lim\limits_{s\rightarrow h(\ell)-}\dfrac{\ell(s)}{\sigma(s)^{2}}$ exist.
@@
    For every $\delta>0$, there exists $\varepsilon>0$ such that
    $\liminf_{n\to\infty}\inf_{\delta n\leq s\leq h(q_{n})-\delta n}{\mathbb{E}}\bigg[(\nu^{n,s}_{1})_{2}\mathbf{1}\bigg\{\nu_{1}^{n,s}\leq\frac{n}{(\log n)^{1+\varepsilon}}\bigg\}\bigg]>0,$ (H2a)
    $\lim_{n\rightarrow\infty}\ \sup_{\delta n\leq s\leq h(q_{n})-\delta n}\frac{\mathbb{E}\big[(\nu_{1}^{n,s})^{3}\big]}{n}=0,,$ (H2b)
    where $(m)_{k}:=m(m-1)\cdots(m-k+1)$ denotes the falling factorial.
Remark.
  • Condition H1(ii) in (H1) ensures that the contribution to the rescaled population size from the very early and very late generations—near the beginning and near extinction—is negligible. Combined with H1(i), this yields the convergence of the rescaled total population size:
    $\frac{1}{n^{2}}\sum_{s\in\mathbb{N}_{+}}q_{n}(s)\;\longrightarrow\;\int_{0}^{\infty}\ell(s)\,\mathrm{d}s.$ (1.1)
@@
    defines a continuous profile. As will be seen in Theorem 1.1, the limiting object involves Brownian motion conditioned on the local time profile $\ell^{\sigma}$. Condition H1(iii) ensures that this conditioning profile has well-defined endpoint values. In many applications, $\sigma_{n}$ is taken to be constant (and hence so is $\sigma$), in which case H1(iii) is automatically satisfied.
  • To clarify the meaning of condition (H2a), note first that uniformly for $\delta n\leq s\leq h(q_{n})-\delta n$,
    ${\mathbb{E}}(\nu^{n,s}_{1})=\frac{q_{n}(s+1)}{q_{n}(s)}=1+o_{n}(1).$
    Thus (H2a) simply requires that the variance $\text{Var}(\nu^{n,s}_{1})$ retain a non‐negligible contribution from the regime $\nu^{n,s}_{1}\leq n/(\log n)^{1+\varepsilon}$. We also record two simple sufficient conditions for (H2a) under the assumption (H1):
    1. (a1)
      There exists $\varepsilon>0$ such that, for every $\delta>0$, uniformly for $\delta n\leq s\leq h(q_{n})-\delta n$,
      ${\mathbb{E}}\big[(\nu^{n,s}_{1})^{3}\big]=o\!\left(\frac{n}{(\log n)^{1+\varepsilon}}\right).$
    2. (a2)
      For every $\delta>0$, the family $\{\nu^{n,s}_{1}:n\geq 1,\ \delta n\leq s\leq h(q_{n})-\delta n\}$ is uniformly $L^{2}$-integrable; that is,
      $\lim_{K\to\infty}\sup_{n\geq 1}\sup_{\delta n\leq s\leq h(q_{n})-\delta n}{\mathbb{E}}\!\left[(\nu^{n,s}_{1})^{2}\mathbf{1}\{\nu^{n,s}_{1}>K\}\right]=0.$
    See Remark 1.3 below for further discussion of conditions (H2a) and (H2b).
@@
Assume (H1)–(H2). Define $\alpha(t):=\frac{1}{4}\int_{0}^{t}\sigma(W^{\ell^{\sigma}}_{s})^{2}\,\mathrm{d}s$. Let $\alpha^{-1}(t)$ be the inverse of $\alpha(t)$. [^2] We will verify in Lemma 2.16 that $\alpha$ is a.s. strictly increasing, so $\alpha^{-1}(t)$ is well-defined. Then
$\Big(n^{-1}C_{\lfloor 2n^{2}t_{1}\rfloor}(\mathscr{T}^{n}),n^{-1}H_{\lfloor n^{2}t_{2}\rfloor}(\mathscr{T}^{n})\Big)_{0\leq t_{1},t_{2}<\infty}\overset{d}{\longrightarrow}\Big(W_{\alpha^{-1}(t_{1})}^{\ell^{\sigma}},W^{\ell^{\sigma}}_{\alpha^{-1}(t_{2})}\Big)_{0\leq t_{1},t_{2}<\infty}$ (1.2)
@@
It is a standard fact that convergence of contour functions implies convergence of the associated trees in the Gromov-Hausdorff topology (see [8, Lemma 2.4]). Applying this, we obtain the following corollary.
@@
Assume (H1)–(H2). Let $\widetilde{\mathscr{T}}^{n}$ be the geometric tree obtained from $\mathscr{T}^{n}$ by assigning length $1/n$ to each edge. Let $\mathcal{T}^{\ell,\sigma}$ denote the compact real tree encoded by the time-changed conditioned Brownian motion $W^{\ell^{\sigma}}_{\alpha^{-1}(t)}$. Then
@@
By Aldous’ theorem [2, Theorem 20], the convergence of the contour functions is equivalent to the combination of tightness and finite-dimensional convergence of the associated tree sequence. In (H2), condition (H2b) is in fact equivalent to the finite-dimensional convergence (see Remark 5.9); while condition (H2a) is closely related to the tightness. In Appendix A, we present a counterexample in which the offspring distribution slightly violates (H2a), leading to a failure of tightness. This indicates a form of sharpness of condition (H2a).
We also remark that the counterexample in Appendix A shows that the condition (H2b), although sufficient for finite-dimensional convergence, is not enough to guarantee tightness.
@@
Our convergence result (1.2) extends to the case where the height profile satisfies $h(q_{n})=h(\ell)=\infty$. In this case, the height and contour processes are defined by following the first infinite branch until it escapes to $\infty$. The associated rescaled processes then converge weakly under the topology of uniform convergence on compact subsets of $\mathbb{R}_{+}$.
The key point is a local dependence property: for both discrete and continuous conditioned (height or contour) functions, the law of the path up to the first hitting time of a height $s$ depends only on the generation profile below height $s$, and coincides with the corresponding law in a finite–height model with the same profile on $[0,s]$. This property is immediate in the discrete case from generation–wise independence, and in the continuous setting, it follows from excursion theory.
This property allows one to define the conditioned process $W^{\ell}$ for profiles $\ell$ with $h(\ell)=\infty$, and together with the finite-height convergence (1.2) and a truncation argument yields the convergence in the infinite-height setting as well.
Remark 1.5.
The convergence results can also be extended to other population models, such as the continuous-time Moran model and the diploid Cannings model (as defined in [15]), using a very similar approach.
@@ 1.2 Related works
A major focus in the study of Cannings models lies in coalescent theory, which describes the genealogies of finite samples via limiting processes such as Kingman’s coalescent [9]. As explained in Remark 5.9, convergence to Kingman’s coalescent is in fact equivalent to the finite-dimensional convergence of the associated tree sequence.
In [14, Theorem 4], Möhle gave the necessary and sufficient condition for convergence to Kingman’s coalescent in the classical Cannings model. Condition (H2b) serves as the analogue of this criterion in the inhomogeneous Cannings setting. Since it is equivalent to the finite-dimensional convergence, it constitutes a necessary condition of the convergence result (1.2), as discussed in Remark 1.3.
@@
It is classical that every continuous bridge function naturally encodes a compact real tree (see Section 2.2.2 for details). For the conditioned Brownian motion $W^{\ell}$, Aldous [3] derived the finite-dimensional distributions of the corresponding encoding tree. Subsequently, Warren [18] provided two alternative, direct proofs of this result. This distribution plays a key role in identifying the limiting process in our setting.
@@
The conditioned Brownian motion $W^{\ell}$ is a self-interacting diffusion constructed by Warren and Yor [20], and independently by Aldous [3]. Detailed constructions will be presented in Section 2.3. Subsequently, Warren [19] further established a close connection between this conditioning problem and a stochastic flow of the Bass-Burdzy type.
More recently, Lupu, Sabot, and Tarrès [12], as well as Aïdékon, Hu, and Shi [1], have further developed this framework by analyzing perturbed reflecting Brownian motions and Brownian loop soups conditioned on local time profiles [^3] Specifically, Lupu, Sabot, and Tarrès [12] studied the Gaussian free field conditioned on local time profile, which is equivalent—via Dynkin’s isomorphism—to a Brownian loop soup of intensity $\frac{1}{2}$, or equivalently, a perturbed reflecting Brownian motion with parameter $2$, conditioned on local time profile. Aïdékon, Hu, and Shi [1] later generalized this result to arbitrary loop soup intensities (or equivalently, arbitrary parameters for perturbed reflecting Brownian motion).. A natural extension of our setting is to incorporate immigration into the inhomogeneous Cannings model. In that case, the rescaled contour and height processes will converge to the (time-changed) perturbed reflecting Brownian motion conditioned on its local time profile.
On the discrete side, the conditional laws of Markov chains and Markovian loop soups given their local time profiles have been studied in [16, 11, 10].
@@ 1.3 Proof outline
Our approach to proving Theorem 1.1 builds on the framework developed by Aldous in [2, Theorem 20], where he showed that for a sequence of random rooted ordered trees, convergence of the contour process is equivalent to the combination of tightness and finite-dimensional convergence of the trees (see Theorem 2.6 for a modified version of this result adapted to our setting). In our case, finite-dimensional convergence is obtained via its equivalence with convergence to Kingman’s coalescent (see Proposition 5.1). The main technical challenge lies in proving tightness. Once tightness is established, the convergence of the contour process follows, and the joint convergence with the height process is then deduced using a deterministic relation between the two processes from [13] (see Section 5.2).
@@
We highlight that the first fact relies on a sequential version of the coming-down-from-infinity (CDFI) property, established in Theorem 3.13. To this end, we introduce a coalescent process, a Markov process that tracks the number of ancestral lineages backward from height $(s+1)\lfloor\delta n\rfloor$. In particular, at height $s\lfloor\delta n\rfloor$ this count is exactly the number of $\delta$-coalescent vertices at the $s$-th height interval. Our assumption (H2a) is used in the proof to control the merge rates of these lineages. Heuristically, under (H2a) the process contracts at a uniform multiplicative rate on mesoscopic windows: on average, over each window of length $\frac{n}{(\log n)^{1+\varepsilon}},$ the lineage count shrinks by a factor $(\log n)^{-c}$ (see Proposition 3.17). So over a time horizon $\delta n$, the cumulative shrinkage is
@@
This results in a rapid decrease in the number of lineages—from $\Theta(n)$ down to a constant—within $\delta n$ steps. By contrast, in the counterexample of Appendix A (where (H2a) fails), a contraction occurs only every $\frac{n}{(\log n)^{1-\varepsilon}}$ steps; over the same horizon $\delta n$ the cumulative shrinkage is
@@
which is insufficient to reduce the lineage count from order $\Theta(n)$ to $O(1)$. This leads to non-convergence of the rescaled contour functions in this example (see Proposition A.1).